The diagonal of a rectangular rug measures 26 ft, and the length is 4 ft more than twice the width. Find the length and width of the rug?
I can't figure this out. My numbers are turning into fractions and I don't think you can factor fractions. Please help.

Answer by msmaeghan(1)

-----------------------------------------------------------------
Hi!
Happy to be of service!
the formula for length is l=4+2w
take the diagonal and make the two pieces into right triangles, and use the Pythagorean theorem.
a^2+b^2=c^2
so now it's:
(4+2w)^2+w^2=26^2
16+4w^2+w^2= 676
5w^2= 660
w^2= 132
w= 11.5
plug the w back in
l= 2(11.5)+4
l= 27
yes, you should get fractions in your answer, and they *are* factorable. so don't worry, you should be fine.